1. Which of the following would not be classified as a current liability in the balance sheet? (a) accounts payable (b) accounts receivable (c) accruals (d) tax payable (e) bank overdraft 2. Which of the following items should be classed as capital expenditure? (a) (b) (c) (d) (e) software maintenance depreciation of equipment research expenses purchase of an automobile interest paid on a loan to purchase inventory 3. Afirm sells goods on credit. The effect will be to (a) (b) (c) (d) (e) decrease inventory, decrease liabilities increase cash, increase liabilities increase cash, decrease inventory decrease inventory, increase other assets decrease inventory, decrease overdraft 4. A company pays cash to a supplier for goods the firm had bought the previous month. The effect will be to: (a) (b) (c) (d) (e) decrease cash, decrease accounts receivable decrease cash, decrease accounts payable decrease cash, increase current liabilities increase cash, decrease cost-of-sales decrease cash, increase cost-of-sales 5. Which of the following would not be classified as a current asset? (a) (b) (c) (d) (e) a tax overpayment security services prepayment work-in-progress precious metal owned by a commodities trader money due to suppliers
6. Omega plc purchased a piece of equipment at a cost of £125,000. It is expected to have a useful economic life of 5 years and a scrap value of £25,000 at the end of its life Omega uses straight line depreciation and a full year of depreciation is charged in the year of purchase What would be the annual depreciation charge and the net book value (NBV) of the equipment at the end of the second year of its life? (a) (b) (c) (d) (e) Depreciation charge £25,000; NBV £100,000 Depreciation charge £25,000; NBV £75,000 Depreciation charge £20,000; NBV £85,000 Depreciation charge £20,000; NBV £80,000 Depreciation charge £30,000; NBV £65,000 7. Using the information given in question 6 above, what would be the depreciation charge in year 2, based on a reducing balance percentage of 30%? (a) £49,000 (b) £73,500 (c) £26,250 (d) £11,250 (e) £37,500 8. The defined benefit pension obligation is not affected by: (a) The rate at which retirement benefits are discounted at (b) How long scheme members are expected to live after they retire () The expected return on the pension funds assets (d) Curent period service costs (e) Changes in scheme terms from final salary to career average 9. The Enterprise value (EV) of a firm is given by (a) (b) (c) (d) (e) Book value of owners equity Market value of owners equity Book value of firms debt plus book value of owners equity Book value of firms debt plus market value of owners equity Market value of firms debt plus market value of owners equity
Answer #1

Ans:

1.(b) account receivable

2. (d) purchase of auto mobile

3. (d) decrease inventory , increase other assets

4.(b) decrease cash, decrease accounts payable

5.(e) money due to suppliers

6. (c) Depreciation 20,000 , NBV = 85,000

Depreciation = 125000 - 25000 / 5  

= 20,000

NBV at second year = 125000 - 20000 -20000

= 85,000

7. (c) 26,250

NBV of asset at year end = 125,000 * 0.7 =  87,500

Depreciation at the of year 2 = 87,500 *0.3

= 26,250

8.

9. (e) Market value of firms debt plus market value of owners equity